I am trying to write an algebraic expreession for the following:
5 less than 2/3 of a number is 3.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

n = a number
(2/3)n - 5 = 3

Answer by stanbon(75887)

5 less than 2/3 of a number is 3.
---
Equation:
(2/3)x-5 = 3
---
(2/3)x = 8
---
(1/3)x = 4
x = 12
=======================
Cheers,
Stan H.